The time T of travel of an automobile trip varies inversely as the speed V. At an average speed of 65mph, a trip took 4 hours. The return trip took 5 hours. Find the average speed of the return trip.
Answer by oberobic(2304)

d = r*t is the basic distance equation.
.
Trip:
r = 65
t = 4
d = 65*4
d = 260
.
Return Trip:
d = 260
t = 5
r = ?
.
t = d/r
t = 260/5
t = 52
.
Done.